You ve collected the following information about St. Pierre, Inc. Sales Net income Dividends Total debt Total equity $130,000 $23,000 $13,000 $74,000 $47,000 Required (a) The sustainable growth rate for St. Pierre, Inc. is 21.28 percent. (Do not include the percent sign (%). Round your answer to 2 decimal places. (e.g., 3216)) (b) If it does grow at this rate, 15747.20 in new borrowing will take place in the coming year, assuming a constant debt-equity ratio. (Do not include the dollar sign (S). Round your answer to 2 decimal places. (e.g., 32.16)) (c) The maximum growth rate that can be supported without any outside financing is percent. (Do not include the percent sign (%). Round your answer to 2 decimal places. (e.g., 3216))
Answer #1

(a) Net Income = $ 23000, Dividends = $ 13000, Total Equity = $ 47000

Retention Ratio = [1-(13000/23000)] ~ 0.435 and Return on Equity (ROE) = (23000 / 47000) x 100 ~ 48.94 %

Sustainable Growth Rate = ROE x Retention Ratio = 48.94 x 0.435 ~ 21.28 %

(b) Current Debt-to-Equity Ratio = 74000 / 47000 ~ 1.574

Debt to Asset Ratio = [1.574 / (1.574+1)] ~ 1.574 / 2.574

Current Asset = 74000 + 47000 = 121000

Expected Assets in Future = Current Asset x (1+ Sustainable Growth Rate) = 121000 x 1.2128 ~ $ 146748.8

New Debt ~ (1.574 / 2.574) x 146748.8 ~ $ 89736.83

New Borrowing = New Debt - Current Debt = 89736.83 - 74000 = $ 15736.83

(c) Maximum Growth Rate that can be supported without any outside financing is the sustainable growth rate calculated in part(a) which is 21.28 %.